Shift report 20000805 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 05 2000 - 07:52:04 EDT

  • Next message: http account: "Shift report 20000805 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Aug  5 07:52:04 2000, Shift: 00:00-08:00
    Supervisor: PS, with: TK
    Log book#3 pages 148 - 152
    
    Shift Summary:
    ------------------------------------------------------------
    1. Cogged beam at gamma = 70 for about 2.5 h.
       Run #2054 has ~820 collision events (trigger 6).
       After run #2054 no beam: "Power Supply Problem".
    2. Run 2054:
       The ZDC(tL-tR) and BB(tL-tR) distributions seem
       to be peaked around 0. 
       Drift time distribution in T3 detector can be found
       in LOG BOOK #3 on page 151.
       After run 2054 we acquired pedestal runs.
    3. HV OK. Nice tracks in MTPC1 and MTPC2.
       Double background peaks in ZDC(tL-tR) distribution.
       TOFW CAMAC readout does not work.      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 05 2000 - 07:55:25 EDT